OpenStudy (anonymous):

Find three points that solve the equation. 2x-3y=-6

OpenStudy (tkhunny):

Well, do it. Just pick three. x = -1 x = 0 x = 1 Find the corresponding value of y for each.

OpenStudy (anonymous):

If I knew how to find the answer I would. I have gotten myself so confused at this point.

OpenStudy (tkhunny):

What are you arguing? Can you substitute x = 0 into the equation you have provided?

OpenStudy (anonymous):

To what I understand, you can substitute it with anything as long as the equation totals the correct number. My problem is finding what those numbers are. Every time I think it's right, I'm told it's wrong.

OpenStudy (tkhunny):

2x-3y=-6 With two variables, you can pick one to be ANYTHING you want. The other comes from the equation. Pick x = 0 2(0) - 3y = -6 0 - 3y = -6 -3y = -6 y = 2 There's a point (0,2) Pick y = 0 2x - 3(0) = -6 2x - 0 = -6 2x = -6 x = -3 There's a point (-3,0) Really, just pick ONE and solve for the other.

OpenStudy (anonymous):

Now THAT makes sense. I will give it another shot. Thank you.